OpenStudy (anonymous):

Rhett decides to build a square room for his movie and music collection. If the area of the room is 9x2 - 6x + 1 square feet, what is the length of one side of the room?

OpenStudy (anonymous):

Choices. (9x - 1) feet (9x + 1) feet (3x + 1) feet (3x - 1) feet

OpenStudy (lgbasallote):

area = s^2 9x^2 - 6x + 1 = s^2 s = \(\sqrt{9x^2 - 6x +1}\) s = \(\sqrt{(3x-1)^2}\) i leave the rest to you

OpenStudy (radar):

It is square. so all sides are equal l=w=x. Area will be the product of two sides. A=9x^2-6x + 1 (3x-1)(3x-1) 3x-1 = length of side.